# AHSG: Adversarial Attacks on High-level Semantics in Graph Neural Networks

### v1

@@ Introduction
$\tau(\alpha_{i,j})=\begin{cases}1,&\text{if }(i=j\text{ and }y_{i}\text{ is unknown})\\ \alpha_{i,j},&\text{if }(y_{i}=y_{j}\text{ and }y_{i}\text{ and }y_{j}\text{ are known})\\ 0,&\text{else}\end{cases}.$ (12)
@@
$\mathrm{\Pi}_{S}(a)=\begin{cases}P_{[0,1]}[a],&\text{if }\sum P_{[0,1]}[a]\leq\epsilon\\ P_{[0,1]}[a-\mu],&\text{if }\mu>0,\sum P_{[0,1]}[a-\mu]=\epsilon\end{cases}.$ (13)
@@
$P_{[0,1]}(x)=\begin{cases}x,&\text{if }x\in[0,1]\\ 1,&\text{if }x>1\\ 0,&\text{if }x<0\end{cases}.$ (14)
Sampling $A$ using $s$ produces the attacked graph $\hat{A}$.The pseudo-code of AHSG is shown in Algorithm 1.
Experiment
In this section, we validate the effectiveness of AHSG through comprehensive experiments. First, we will introduce the experimental setup, including dataset statistics, baselines, GNNs, and parameter settings. Next, we present the experimental results and analyze them to verify the performance and characteristics of AHSG.
Dataset
We evaluated AHSG on three well-known datasets: Cora, Citeseer, and Cora-ML. All datasets contain unweighted edges, allowing the generation of an adjacency matrix $A$, and include sparse bag-of-words feature vectors that can be used as input for GCN. Among them, Cora and Citeseer are symmetric matrices (i.e., undirected graphs), while Cora-ML is an asymmetric matrix (i.e., directed graph). For this experiment, we converted the Cora-ML dataset into an undirected graph. The Cora and Citeseer datasets have binary features, while the Cora-ML dataset contains continuous features. Following the setup from previous work (Kipf and Welling 2017), the feature vectors of all nodes are fed into the GCN, with only 140 and 120 training nodes for Cora and Citeseer, respectively, and 200 training nodes for the Cora-ML dataset. The number of test nodes is 1,000 for all three datasets. The statistical results are shown in Table 1.
@@
Algorithm 1 AHSG
@@
Parameter: $L$,$\lambda$,$\beta$,$T$,$\eta_{t}$,$K$,$h$
@@
1:  Train GCN to obtain $W_{(1,2)}^{\ast}$.
2:  Calculate $H_{1}$ via Equation(4).
3:  for $l\leftarrow 1\text{ to }L$ do
4:     Calculate $\widehat{H_{1}}$ via Equation(9).
5:     Gradient descent:$\alpha(l)=\alpha(l-1)-\lambda\nabla\text{loss}_{1}\left(\alpha(l-1)\right)$.
@@
9:     Gradient descent:$a(t)=s(t-1)-\eta_{t}\nabla\text{loss}_{2}(s(t-1))$
@@
Baselines and GNNs
The baselines include Random and DICE (Waniek et al. 2018), which are based on random sampling, as well as Meta-Self (Zügner and Günnemann 2019), PGD (Xu et al. 2019), and GradArgmax (Dai et al. 2018). Meta-Self employs meta-learning to attack node classification models. PGD proposes a topology attack framework based on edge perturbation, which overcomes the challenge of attacking discrete graph structures from a first-order optimization perspective. GradArgmax modifies one edge in the current graph at a time using a gradient-based greedy algorithm, and iterates until the attack budget is exhausted.
We use AHSG and the aforementioned baselines to attack GCN and five defense or robust GNN models. Jaccard (Wu et al. 2019) calculates the Jaccard similarity of connected node pairs and retains only those links with high similarity. Svd (Entezari et al. 2020) utilizes low-order approximations of the graph to enhance the performance of GCN against adversarial attacks. ProGNN (Jin et al. 2020) jointly learns the graph structure and a robust GNN model from the perturbed graph. The SimPGCN (Jin et al. 2021b) framework enhances GCN robustness by effectively preserving node similarity. RGCN (Zhu et al. 2019) proposes that latent representations based on Gaussian distributions can effectively absorb the impact of adversarial attacks.
Experimental Setup
We conduct comparative experiments with advanced graph attack and defense methods. For Svd and Jaccard, we adjust the order of high-order approximations and the threshold for removing low-similarity links to achieve optimal defense performance. Other attack methods use default parameters. In our method, the default settings are as follows: the number of iterations $L$ for perturbation is set to 300, with a learning rate $\lambda$ of 0.3. The regularization coefficient $\beta$ is 0.2. For reconstruction, the number of iterations $T$ is 300, and the dynamic learning rate $\eta_{t}$ is $\frac{q}{\sqrt{t+1}}$, where $t$ denotes the iteration count and $q$ is set to 20. The number of selections $K$ for the probability matrix is 20. When using a two-layer GCN as the victim model, the dimension $h$ of the first layer is 128, and the second layer’s dimension corresponds to the number of classes. In CSBMs, the number of generated nodes $n$ is 1000. Node feature mean $\mu$ is calculated as $\frac{M\sigma}{2\sqrt{d}}$, with variance $\sigma=1$, $d=\frac{n}{(\ln(n))^{2}}=21$, and $M=0.5$. Smaller value of $M$ causes the reference classifier to rely more on structural information. Since AHSG focuses on structural attacks, it is essential for the reference classifier to analyze the semantics within the structure. The edge probability $p$ between same-class nodes is 0.6326%, while the edge probability $q$ between different-class nodes is 0.1481%.
Attack performance on GNNs
@@
Except for ProGNN, other GNNs use the evasion attack setting. ProGNN performs both graph structure purification and parameter learning simultaneously, so the attack occurs before training. The attack results on different GNNs by modifying 10% of the edges are shown in Table 2. We observe that all baseline methods lead to a performance drop in the victim models. AHSG outperforms all other attackers on all datasets. Furthermore, AHSG achieves the best attack performance on GCN among different GNN models, as GCN is the corresponding surrogate model, indicating a white-box attack scenario. Additionally, even against the five defensive GNNs, AHSG still demonstrates strong attack performance, highlighting its robust generalization capability derived from the surrogate GCN model.
@@
Dataset Method Clean Random DICE Meta-Self GradArgmax PGD AHSG
Cora GCN 0.8230 0.8090 0.7970 0.8090 0.7160 0.7080 0.6460
Jaccard 0.7860 0.7840 0.7630 0.7360 0.7120 0.7590 0.6760
Svd 0.7290 0.7110 0.7070 0.6650 0.6730 0.6860 0.6480
ProGNN 0.8090 0.7940 0.7740 0.6590 0.7370 0.7030 0.6540
SimPGCN 0.7900 0.7820 0.7770 0.7840 0.7030 0.7280 0.6820
RGCN 0.8010 0.7480 0.7450 0.7530 0.6900 0.6780 0.6510
Citeseer GCN 0.6660 0.6550 0.6470 0.6500 0.5780 0.5930 0.5280
Jaccard 0.6650 0.6320 0.6490 0.6450 0.6170 0.6230 0.5760
Svd 0.6010 0.5930 0.5740 0.5730 0.5880 0.5650 0.5580
ProGNN 0.6830 0.6250 0.6250 0.5310 0.5660 0.6030 0.5300
SimPGCN 0.6560 0.6460 0.6500 0.6480 0.5850 0.6150 0.5810
RGCN 0.6100 0.5790 0.5800 0.5690 0.5760 0.5730 0.5340
Cora-ML GCN 0.8590 0.8510 0.8490 0.8270 0.7430 0.7910 0.7330
Jaccard 0.8600 0.8390 0.8360 0.8070 0.7700 0.8280 0.7620
Svd 0.8200 0.7990 0.8260 0.7940 0.7590 0.8160 0.7570
ProGNN 0.8400 0.8050 0.8050 0.7910 0.7490 0.8230 0.7350
SimPGCN 0.8460 0.8360 0.8390 0.8170 0.7430 0.8180 0.7400
RGCN 0.8590 0.8420 0.8420 0.8310 0.7540 0.8280 0.7550
@@
Dataset Attack Ratio Random DICE Meta-Self GradArgmax PGD AHSG
Cora 5% 0.814 0.805 0.814 0.765 0.756 0.698
10% 0.809 0.797 0.809 0.716 0.708 0.646
15% 0.795 0.777 0.805 0.685 0.681 0.615
20% 0.787 0.771 0.798 0.652 0.645 0.585
Citeseer 5% 0.657 0.659 0.651 0.608 0.610 0.565
10% 0.655 0.647 0.650 0.578 0.593 0.528
15% 0.653 0.642 0.631 0.539 0.564 0.495
20% 0.641 0.634 0.628 0.507 0.529 0.487
Cora-ML 5% 0.853 0.850 0.841 0.791 0.843 0.776
10% 0.851 0.849 0.827 0.743 0.791 0.733
15% 0.846 0.843 0.821 0.721 0.751 0.700
20% 0.837 0.824 0.809 0.695 0.708 0.685
Table 3: Accuracy of GCN under different attack budgets. The best result in each row is highlighted in bold.
Attack Performance w.r.t attack budget
@@
Semantic Detection
In the fields of computer vision (CV) and natural language processing (NLP), humans can effectively perform semantic checks. However, in the domain of graph, it is challenging for humans to inspect the semantics of large-scale graphs. To address this, (Gosch et al. 2023) attempts to define semantic boundaries by introducing a reference classifier $g$ to represent changes in semantic content. The reference classifier $g$ can be derived from knowledge about the data generation process. According to the data generation process of CSBMs, (Gosch et al. 2023) uses a Bayes classifier as $g$.
@@
$G(X^{\prime},A^{\prime})=\text{attack}(G(X,A)).$ (16)
$g(X,A)_{v}=\arg\max_{y}\left(Bayes(X,A,v,y)\right).$ (17)
$Bayes(X,A,v,y)=log(p\left(X_{v}\middle|y_{v}\right)+$ (18)
@@
We reclassify the nodes in the attacked graph $G(X^{\prime},A^{\prime})$, which is generated by Equation (16). If the results satisfy (15), we consider that the attack method does not significantly disrupt the semantics of node $v$ in the graph. The number of nodes with consistent classification results before and after the attack can be regarded as the degree of graph semantics preservation. We denote this by Bayes_maintain, as shown in Equation (17), where $\mathbb{I}$ is an indicator function. As shown in Table 4, the Bayes reference classifier achieves 92.4% accuracy on clean graphs, validating its effectiveness as a semantic proxy. Comparison with other attack methods reveals that AHSG alters the Bayesian classification results the least, thereby maintaining the majority of the primary semantics in the classification task while achieving the best attack performance.
@@
Related Work
In graph modification attack, (Zügner, Akbarnejad, and Günnemann 2018) introduced the first adversarial attack method on graph, called Nettack. (Zügner and Günnemann 2019) treated the input graph as a hyperparameter to be learned and modified one edge per iteration. (Xu et al. 2019) overcame the challenge of attacking discrete graph structure data. (Dai et al. 2018) employed reinforcement learning, gradient-based greedy algorithms, and genetic algorithms to attack GNNs in various scenarios. (Liu et al. 2022) generated unweighted gradients on the graph structure unaffected by node confidence, fully utilizing the attack budget. However, the issue of altering graph semantics during an attack has not been well addressed in previous work. Although the attack budget is minimal relative to the entire graph, it can still lead to significant changes in the semantics of the attacked graph. Some works maintained certain properties of the graph to achieve imperceptibility, such as Nettack (Zügner, Akbarnejad, and Günnemann 2018) and HAO (Chen et al. 2022). But these constraints are necessary but not sufficient conditions for preserving semantics.
Conclusion
In this work, we study graph structure attack on GNNs under the evasion attack setting. Unlike other studies that regard attack budgets or graph properties as semantic proxies, we propose AHSG, which preserves the primary semantics of the GNN’s hidden layers to maintain the overall semantics of the graph, thereby eliminating the potential for significant semantic disruption under insufficient constraints. Specifically, by leveraging the similarity of primary semantics among same-class nodes, we construct specific representations that lead to fail of GNNs while preserving primary semantics within the set of same-class nodes. For the generation of adversarial examples, we employ the PGD algorithm to reconstruct the graph structure from the perturbed latent representations. Additionally, a regularization term is incorporated during the latent representations perturbation to further constrain semantic invariance by maintaining similarity to the original representations. Extensive experimental results on different datasets demonstrate the superior attack performance of the proposed AHSG on various types of GNNs, including GCN and multiple defense models. Notably, semantic detection of the attacked graph indicates that AHSG largely preserves the primary semantics relevant to the current task.
Appendix A Appendix
Ablation Study
@@ Appendix A Appendix
Clean 0.823 0.666 0.8590
AHSG-rec 0.8090 0.6550 0.8520
AHSG-hid 0.804 0.650 0.8420
AHSG 0.698 0.565 0.7760

### v2

@@ 4.2 Model Optimization
$\tau(\alpha_{i,j})=\begin{cases}1,&\text{if }(i=j\text{ and }y_{i}\text{ is unknown})\\ \alpha_{i,j},&\text{if }(y_{i}=y_{j}\text{ and }y_{i}\text{ and }y_{j}\text{ are known})\\ 0,&\text{else}.\end{cases}$ (12)
@@
$\mathrm{\Pi}_{S}(a)=\begin{cases}P_{[0,1]}[a],&\text{if }\sum P_{[0,1]}[a]\leq\epsilon\\ P_{[0,1]}[a-\mu],&\text{if }\mu>0,\sum P_{[0,1]}[a-\mu]=\epsilon,\end{cases}$ (13)
@@
$P_{[0,1]}(x)=\begin{cases}x,&\text{if }x\in[0,1]\\ 1,&\text{if }x>1\\ 0,&\text{if }x<0.\end{cases}$ (14)
Sampling $A$ using $s$ produces the attacked graph $\hat{A}$. The overall procedure of AHSG is shown in Algorithm 1.
5 Experiment
In this section, we validate the effectiveness of AHSG through comprehensive experiments.
5.1 Experimental Settings
We evaluated AHSG on three well-known datasets: Cora, Citeseer, and Cora-ML. The statistical results are shown in Table 1. All datasets contain unweighted edges, allowing the generation of an adjacency matrix $A$, and include sparse bag-of-words feature vectors that can be used as input for GNN. Among them, Cora and Citeseer are symmetric matrices (i.e., undirected graphs), while Cora-ML is an asymmetric matrix (i.e., directed graph). For this experiment, we converted the Cora-ML dataset into an undirected graph. The Cora and Citeseer datasets have binary features, while the Cora-ML dataset contains continuous features. Following the setup from previous work, the feature vectors of all nodes are fed into the GNN, with only 140 and 120 training nodes for Cora and Citeseer, respectively, and 200 training nodes for the Cora-ML dataset. The number of test nodes is 1,000 for all three datasets.
The baselines include Random, Meta-Self Zügner and Günnemann (2019), GradArgmax Dai et al. (2018), PGD Xu et al. (2019), EpoAtk Lin et al. (2020) and GraD Liu et al. (2022). We choose GCN as the surrogate model, where the total number of layers $L$ in the neural network is 2, and $l$ is 1. We conduct comparative experiments with advanced graph attack and defense methods. For Svd and Jaccard, we adjust the order of high-order approximations and the threshold for removing low-similarity links to achieve optimal defense performance. Other attack methods use default parameters.
@@ 5.1 Experimental Settings
In our method, the default settings are as follows: the number of iterations $E$ for perturbation is set to 300, with a learning rate $\lambda$ of 0.3. The regularization coefficient $\beta$ is 0.2. For reconstruction, the number of iterations $T$ is 300, and the dynamic learning rate $\eta_{t}$ is $\frac{q}{\sqrt{t+1}}$, where $t$ denotes the iteration count and $q$ is set to 20. The number of selections $K$ for the probability matrix is 20. When using a two-layer GCN as the victim model, the dimension $h$ of the first layer is 128, and the second layer’s dimension corresponds to the number of classes.
Algorithm 1 The procedure of AHSG
@@
Parameter: $E$,$\lambda$,$\beta$,$T$,$\eta_{t}$,$K$,$h$
@@
1:  Train GNN to obtain $W_{(1,2)}^{\ast}$.
2:  Calculate $H_{l}$ via Equation(4).
3:  for $e\leftarrow 1\text{ to }E$ do
4:     Calculate $\widehat{H_{l}}$ via Equation(9).
5:     Gradient descent:$\alpha(e)=\alpha(e-1)-\lambda\nabla\mathcal{L}_{1}\left(\alpha(e-1)\right)$.
@@
9:     Gradient descent:$a(t)=s(t-1)-\eta_{t}\nabla\mathcal{L}_{2}(s(t-1))$
@@
5.2 Attack Performance on GNNs
@@
Except for ProGNN, other GNNs use the evasion attack setting. ProGNN performs both graph structure purification and parameter learning simultaneously, so the attack occurs before training. The attack results on different GNNs by modifying 10% of the edges are shown in Table 2. We observe that all baseline methods lead to a performance drop in the victim models. AHSG outperforms all other attackers on all datasets. Furthermore, AHSG achieves the best attack performance on GCN among different GNN models, since GCN is the corresponding surrogate model, indicating a white-box attack scenario. Furthermore, even against the five defensive GNNs, AHSG still demonstrates strong attack performance, highlighting its robust generalization capability derived from the surrogate model.
@@
Dataset Method Clean Random Meta-Self GradArgmax PGD EpoAtk GraD AHSG
Cora GCN 0.823 0.809 0.809 0.716 0.708 0.685 0.673 0.646
Jaccard 0.786 0.784 0.736 0.712 0.759 0.693 0.703 0.676
SVD 0.729 0.711 0.665 0.673 0.686 0.689 0.715 0.648
ProGNN 0.809 0.794 0.659 0.737 0.703 0.673 0.665 0.654
SimPGCN 0.790 0.782 0.784 0.703 0.728 0.703 0.695 0.682
RGCN 0.801 0.748 0.753 0.690 0.678 0.671 0.664 0.651
Citeseer GCN 0.666 0.655 0.650 0.578 0.593 0.549 0.542 0.528
Jaccard 0.665 0.632 0.645 0.617 0.623 0.591 0.622 0.576
SVD 0.601 0.593 0.573 0.588 0.565 0.589 0.574 0.558
ProGNN 0.683 0.625 0.531 0.566 0.603 0.597 0.582 0.530
SimPGCN 0.656 0.646 0.648 0.585 0.615 0.613 0.602 0.581
RGCN 0.610 0.579 0.569 0.576 0.573 0.559 0.552 0.534
Cora-ML GCN 0.859 0.851 0.827 0.743 0.791 0.741 0.738 0.733
Jaccard 0.860 0.839 0.807 0.770 0.828 0.801 0.813 0.762
SVD 0.820 0.799 0.794 0.759 0.816 0.763 0.808 0.757
ProGNN 0.840 0.805 0.791 0.749 0.823 0.783 0.769 0.735
SimPGCN 0.846 0.836 0.817 0.743 0.818 0.774 0.758 0.740
RGCN 0.859 0.842 0.831 0.754 0.828 0.785 0.771 0.755
@@
Dataset Attack Ratio Random Meta-Self GradArgmax PGD EpoAtk GraD AHSG
Cora 5% 0.814 0.814 0.765 0.756 0.731 0.725 0.698
10% 0.809 0.809 0.716 0.708 0.685 0.673 0.646
15% 0.795 0.805 0.685 0.681 0.634 OOM 0.615
20% 0.787 0.798 0.652 0.645 0.607 OOM 0.585
Citeseer 5% 0.657 0.651 0.608 0.610 0.585 0.578 0.565
10% 0.655 0.650 0.578 0.593 0.549 0.542 0.528
15% 0.653 0.631 0.539 0.564 0.523 OOM 0.495
20% 0.641 0.628 0.507 0.529 0.501 OOM 0.487
Cora-ML 5% 0.853 0.841 0.791 0.843 0.795 0.789 0.776
10% 0.851 0.827 0.743 0.791 0.741 0.738 0.733
15% 0.846 0.821 0.721 0.751 0.711 OOM 0.700
20% 0.837 0.809 0.695 0.708 0.698 OOM 0.685
Table 3: Accuracy of GCN under different attack budgets. The best result in each row is highlighted in bold. OOM represents Out of Memory.
5.3 Attack Performance w.r.t Attack Budget
@@
5.4 Semantic Detection
In the fields of computer vision (CV) and natural language processing (NLP), humans can effectively perform semantic checks. However, in the domain of graph, it is challenging for humans to inspect the semantics of large-scale graphs. To address this, Gosch et al. (2023) attempt to define semantic boundaries by introducing a reference classifier $g$ to represent changes in semantic content. The reference classifier $g$ can be derived from knowledge about the data generation process. According to the data generation process of CSBMs, Gosch et al. (2023) use a Bayes classifier as $g$.
@@ 5.4 Semantic Detection
We reclassify the nodes in the attacked graph $G(X^{\prime},A^{\prime})$, which is generated by Equation (15).
$G(X^{\prime},A^{\prime})=\text{attack}(G(X,A)).$ (15)
As shown in Equation (16), if the reference classifier $g$ gives the same prediction for node $v$ in both the original and attacked graph, and the prediction is correct, then Gosch et al. (2023) consider that the primary semantics of node $v$ have not been altered by the attack.
@@
In Equation (17), the reference classifier $g$ consists of a feature-based Bayesian classifier and a structure-based Bayesian classifier. The structure-based Bayesian classifier considers four scenarios: whether nodes of the same class are connected and whether nodes of different classes are connected.
$g(X,A)_{v}=\arg\max_{y}\left(Bayes(X,A,v,y)\right),$ (17)
$Bayes(X,A,v,y)=log(p\left(X_{v}\middle|y_{v}\right)+$
@@
The number of nodes with consistent classification results before and after the attack can be regarded as the degree of graph semantics preservation. We denote this by Bayes_maintain, as shown in Equation (18), where $\mathbb{I}$ is an indicator function.
@@
As shown in Table 4, the Bayes reference classifier achieves 92.4% accuracy on clean graphs, validating its effectiveness as a semantic proxy. Comparison with other attack methods reveals that AHSG alters the Bayesian classification results the least, thereby maintaining the majority of the primary semantics in the classification task while achieving the best attack performance.
@@
Note that in CSBMs, We use the original default settings. The number of generated nodes $n$ is 1000. Node feature mean $\mu$ is calculated as $\frac{M\sigma}{2\sqrt{d}}$, with variance $\sigma=1$, $d=\frac{n}{(\ln(n))^{2}}=21$, and $M=0.5$. Smaller value of $M$ causes the reference classifier to rely more on structural information. Since AHSG focuses on structural attacks, it is essential for the reference classifier to analyze the semantics within the structure. The edge probability $p$ between same-class nodes is 0.6326%, while the edge probability $q$ between different-class nodes is 0.1481%.
5.5 Ablation Study
@@
Clean 0.823 0.666 0.859
AHSG-rec 0.809 0.655 0.852
AHSG-hid 0.804 0.650 0.842
AHSG 0.698 0.565 0.776
@@ 5.6 Parameter Analysis
6 Conclusion
We investigate graph structure attacks on GNNs under the evasion setting and propose AHSG, which preserves primary semantics in hidden layers to prevent significant semantic disruption. By exploiting the similarity of primary semantics among nodes with the same class, AHSG constructs representations that cause GNN failure while maintaining semantic consistency. Adversarial graphs are generated using the PGD algorithm, with a regularization term added to ensure semantic invariance. Experiments on various datasets show AHSG’s superior attack performance against multiple GNNs, including defense models. Semantic detection confirms that AHSG effectively preserves task-relevant primary semantics.